A solution is made containing 14.8 g of CH3OH in 186 g H2O.

a. Calculate the mole fraction of CH3OH.
b. Calculate the mass percent of CH3OH.
c. Calculate the molality of CH3OH. (mol/kg)
Answer #1

Answer:-

Given:-

weight or mass of CH3OH (mCH3OH) = 14.8 g

weight or mass of H2O (mH2O) = 186 g

As we know that

molar mass of  CH3OH = 32 g mol-1

molar mass of H2O = 18 g mol-1

(a)-

As we know that

no. of moles of compound = wt. or mass of compound / molar mass of compound

therefore

no. of moles of CH3OH (nCH3OH) = weight or mass of CH3OH (mCH3OH) / molar mass of CH3OH

no. of moles of CH3OH (nCH3OH) = 14.8 g / 32 g mol-1

no. of moles of CH3OH (nCH3OH) = 0.4625 mol

similarly

no. of moles of H2O (nH2O) = weight or mass of H2O (mH2O) / molar mass of H2O

no. of moles of H2O (nH2O) = 186 g / 18 g mol-1

no. of moles of H2O (nH2O) = 10.33 mol

so

total no. of moles present in solution (ntotal) = no. of moles of H2O (nH2O) + no. of moles of CH3OH (nCH3OH)

total no. of moles present in solution (ntotal) = 10.33 mol + 0.4625 mol

total no. of moles present in solution (ntotal) = 10.7925 mol

also we know that

mole fraction of CH3OH = no. of moles of CH3OH (nCH3OH) / total no. of moles present in solution (ntotal)

mole fraction of CH3OH = 0.4625 mol / 10.7925 mol

mole fraction of CH3OH = 0.04285 (i.e the answer)

(b)-

As we know that

total mass of solution (mtotal) = weight or mass of H2O (mH2O) + weight or mass of CH3OH (mCH3OH)

total mass of solution (mtotal) =  186 g + 14.8 g

total mass of solution (mtotal) = 200.8 g

Also we know that

mass percent of CH3OH (%) = weight or mass of CH3OH (mCH3OH) 100 / total mass of solution (mtotal)

mass percent of CH3OH (%) = 14.8 g 100 / 200.8 g

mass percent of CH3OH (%) = 1480 / 200.8

mass percent of CH3OH (%) = 7.371 % ( i.e the answer)

(c)-

As we know that

weight or mass of CH3OH (mCH3OH) = 14.8 g

weight or mass of H2O (mH2O) = 186 g = 186 / 1000 = 0.186 Kg

As we know that

molar mass of  CH3OH = 32 g mol-1

As we know that

molality of the compound = wt. of compound / molar mass of compound wt of solvent in Kg

molality of the CH3OH = wt. of CH3OH / molar mass of CH3OH wt of H2O in Kg

molality of the CH3OH = 14.8 g / 32 ​​​​​​​ 0.186 Kg

molality of the CH3OH = 0.4625 mol / 0.186 Kg

molality of the CH3OH = 2.4866 mol/ Kg (i.e the answer)